OpenStudy (anonymous):

A geologist calculated the following depths of various locations of a large lake, relative to sea level: -20, 6, -4, 12, 17, 33, -14, -2. Put the depths in order and select the median from the choices below. Answer 2 -4 -7 15

OpenStudy (anonymous):

@Aperogalics

OpenStudy (anonymous):

-20,-14,-4,-2,6,12,17,33 so 6-2=4/2=2:)